OpenStudy (anonymous):

under the QUADRATIC EQUATION(Roots) Find two numbers whose sum is 28 and whose product is 96.

OpenStudy (saifoo.khan):

from how many years you arent doing your hw? just wondering..

OpenStudy (anonymous):

just now. Because I totally not understand what my Prof teaches us

OpenStudy (luffingsails):

Consider what you know from what is given. Two numbers, X and Y. Given: X + Y = 28 and X*Y = 96. Pick one of the givens, solve for X or Y and then substitute into the other equation. That will get you started.

OpenStudy (saifoo.khan):

xy = 96 x + y = 28

OpenStudy (saifoo.khan):

substituing the eq1 into eq2 x = 96/y x + y = 28 96/y + y = 28 96 + y^2 =28y y^2 -28y +96 = 0 y = 24, y = 4

OpenStudy (saifoo.khan):

now substitute them in the eq1 to get 2 values